# Exact predicates, exact constructions and combinatorics

### v1

@@ 3.2. ThingiCSG
  • the files from my non-regression test suite, with small but challenging examples, with degeneracies, as well as larger ones, such has the “Finobunny shere” on the first page.
@@
Both OpenSCAD files and the .csg parser will be made available in a new thinggiCSG repository, to make it easy to test and benchmark new research projects on mesh CSG.
@@
  • the algorrithm presented in this article.
@@ 4. Conclusions
This work can be improved and extended in several directions. Clearly, the main missing component is a “snap rounding” algorithm, that transforms the exact coordinates into floating point numbers while preserving some topological properties. The approach described in (Devillers et al., 2018; Valque, 2019) is very promising. Another interesting direction is the approach completely based on floating-point numbers described in (Smith and Dodgson, 2006; Lalish, 2022). It is a completely different paradigm. While the “manifold” OpenSCAD kernel, based on this paradigm, does not always output a correct result, it does very often, at a spectacular speed. It is very rare that it takes more than 1s for a CSG. Is it possible to derive an algorithm with the same performance and more guarantees ? Finally, it may be interesting to enrich the ThingiCSG database introduced in this article with a larger set of examples, especially if a larger research community wants to tackle this type of problems.
Acknowledgements
This work is supported by the COSMOGRAM Inria exploratory action (AeX). I wish to thank Nicolas Ray for discussions, and for always encouraging me to explore the simplest solution first, such as the global vertex map.

### v2

@@ 3.3. ThingiCSG
  • the files from my non-regression test suite, with small but challenging examples, with degeneracies, as well as larger ones, such has the “Fibobunny shere” on the first page.
@@
Both OpenSCAD files and the .csg parser are available in a new thingiCSG repository, to make it easy to test and benchmark new research projects on mesh CSG.
@@
  • the algorithm presented in this article.
@@ 4. Conclusions
This work can be improved and extended in several directions. Clearly, the main missing component is a “snap rounding” algorithm, that transforms the exact coordinates into floating point numbers while preserving some topological properties. The approach described in (Devillers et al., 2018; Valque, 2019) is very promising. Another interesting direction is the approach completely based on floating-point numbers described in (Smith and Dodgson, 2006; Lalish, 2022). It is a completely different paradigm. While the “manifold” OpenSCAD kernel, based on this paradigm, does not always output a correct result, it does very often, at a spectacular speed. It is very rare that it takes more than 1s for a CSG. Is it possible to derive an algorithm with the same performance and more guarantees? It is also important to mention here that if using integer-only coordinates is allowed by the application context, a completely different class of methods can be used, such as the ones presented in (Trettner et al., 2022; Nehring-Wirxel et al., 2021). They result in a spectacular acceleration factor (50x to 100x) as compared to what is presented here. Finally, it may be interesting to enrich the ThingiCSG database introduced in this article with a larger set of examples, especially if a larger research community wants to tackle this type of problems.
Acknowledgements.
I wish to thank Nicolas Ray for many discussions and for vigorously encouraging me to explore the simplest option first. This research was supported by the Inria AeX COSMOGRAM-Launchpad grant.